Custom Query (432 matches)


Show under each result:

Results (346 - 348 of 432)

Ticket Resolution Summary Owner Reporter
#354 fixed Hawk 2010 spectral calibration shifted mggr mggr

On examination, the Hawk spectral calibration appears to have shifted between 2009 and 2010. However, when processing using the new, shifted calibration, we see absorption features misplaced by the same amount as the shift.

This implies the spectral calibration is incorrect.

#34 fixed 138-jitter ATM sbg amro

The ATM lines have jitter of 2 pixels. attached screenshot shows that. Steve needs to make decision about data delivery as it is ready to be burnt and shipped.

#101 fixed Support: 22/Jan/2007, Mohammed Haq (Mark Cutler), GB04/29 sbg mggr

Contacted by Bill handing over Mohammed's processing problem. He was having difficulties with DEMs (specifically formatting.

Note: See TracQuery for help on using queries.